Your toss a ball straight up with an initial speed of 6.4 m s-1. Neglecting air resistance, how long is it in the air above your height?

Answer seconds
Answer #1

the equation for time of flight is

T = time of asent + time of desent

= u /g + u /g

= 2u /g

= 2 (6.4 ) /9.8

= 12.8 / 9.8

T = 1.306 sec